OpenStudy (anonymous):

simplify x-2 over x3-8

OpenStudy (anonymous):

\[x^3-8=(x-2)(x^2+2x+4)\]

OpenStudy (anonymous):

cancel to get \[\frac{x-2}{x^3-8}=\frac{1}{x^2+2x+4}\]